# Robust $M$-Estimation of Scatter Matrices via Precision Structure Shrinkage

### v1

@@ 6 Proofs
Consider the sequence $\{\mathrm{Y}_{\ell}\}_{\ell\in\mathbb{N}}$ that satisfies $\mathrm{Z}_{\ell}=\mathrm{X}\cup\mathrm{Y}_{\ell}\in\mathcal{E}_{m}(X)$ for $\ell\in\mathbb{N}$. Let
@@
By the compactness of $\mathcal{S}_{++}^{p*}\coloneqq\{\mathrm{X}\in\mathcal{S}_{++}^{p};\,\mathrm{Tr}[X]=1\}$ in the Euclidean topology, the sequence $\{\Omega_{\ell}\}_{\ell\in\mathbb{N}}$ admits a convergent subsequence $\{\Omega_{j}\}_{j\in J}$. Hereafter, we shall denote this limit operation by $\Omega_{j}\xrightarrow{j\in J}\Omega\in\mathcal{S}_{+}^{p*}$ or
$\lim_{j\in J}\Omega_{j}=\Omega\in\mathcal{S}_{+}^{p*}.$
@@
for $i\in[m]$. Since $\mathbb{S}^{p-1}$ is compact in the Euclidean topology, the product space $\mathcal{S}_{++}^{p*}\times\mathbb{S}^{p-1}$ is also compact in the product topology, and hence there exists an index set $J$ and a point $\bm{u}_{i}\in\mathbb{S}^{p-1}$ such that $\bm{u}_{i,j}\xrightarrow{j\in J}\bm{u}_{i}$. Let $\bar{P}_{\Omega}\coloneqq I_{p}-\Omega\Omega^{\dagger}$ denote the orthogonal projection onto $\textnormal{Ker}(\Omega)$. It holds that
@@
Hence, (18) is a necessary condition for $\mathrm{Z}\in\mathcal{E}_{m}(X)$. If $\Omega\bm{x}_{i}\neq\bm{0}$, then
@@
$\psi(\bm{x}_{i}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{x}_{i})\frac{\bm{x}_{i}^{\top}\Omega_{j}^{1/2}\bar{P}_{\Omega}\Omega_{j}^{1/2}\bm{x}_{i}}{\bm{x}_{i}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{x}_{i}}\xrightarrow{j\in J}0.$
@@
$\psi(\bm{y}_{i,j}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{y}_{i,j})\frac{\bm{u}_{i,j}^{\top}\Omega_{j}^{1/2}\bar{P}_{\Omega}\Omega_{j}^{1/2}\bm{u}_{i,j}}{\bm{u}_{i,j}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{u}_{i,j}}\xrightarrow{j\in J}0.$
@@
$\textnormal{rank}(\bar{P}_{\Omega})=\lim_{j\in J}\frac{1}{n+m}\sum_{\bm{z}\in\mathrm{Z}_{0,j}}\psi(\bm{z}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{z})\frac{\bm{z}^{\top}\Omega_{j}^{1/2}\bar{P}_{\Omega}\Omega_{j}^{1/2}\bm{z}}{\bm{z}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{z}},$ (19)
@@
$p$ $\geq\lim_{j\in J}\frac{1}{n+m}\left\{\sum_{\bm{z}\in\mathrm{Z}_{0,j}}\psi\left(\bm{z}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{z}\right)\frac{\bm{z}^{\top}\Omega_{j}\bm{z}}{\bm{z}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{z}}\right$.
@@
$\geq\lim_{j\in J}\frac{1}{n+m}\left\{\sum_{\bm{z}\in\mathrm{Z}_{0,j}}\psi\left(\bm{z}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{z}\right)\frac{\bm{z}^{\top}\Omega_{j}^{1/2}\bar{P}_{\Omega}\Omega_{j}^{1/2}\bm{z}}{\bm{z}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{z}}\right$.
@@
$p$ $\geq\textnormal{rank}(\bar{P}_{\Omega})+\frac{1}{n+m}\sum_{\bm{x}_{i}\in\mathrm{X}\cap\textnormal{Ker}(\Omega)^{c}}\lim_{j\in J}\psi\left(\bm{x}_{i}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{x}_{i}\right)\frac{\bm{x}_{i}^{\top}\Omega_{j}\bm{x}_{i}}{\bm{x}_{i}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{x}_{i}}.$
Letting $\omega_{j}\xrightarrow{j\in J}+\infty$ yields
@@
When $\varepsilon_{m}>1-p/\kappa$, setting $Y=\{\bm{0},\ldots,\bm{0}\}$ yields $\hat{P}_{n+m}(\{\bm{0}\})>1-p/\kappa$, which in turn implies that $Z\notin\mathcal{E}_{m}(X)$, and hence $\hat{V}(Z)$ breaks down, in accordance with Lemma 4.1-(2). When $\varepsilon_{m}<1-p/\kappa-(p-1)/n$, from Lemmas 4.1-(1) and 4.2, we see that $\hat{V}(Z)$ does not break down. ∎
@@ 7 Concluding remarks
We further examined statistical properties of the proposed method from the perspectives of penalized $M$-estimation and Bayesian estimation. Moreover, as summarized in Figure 1, we discussed correspondences between the proposed estimator and various existing estimators. Through numerical experiments, we demonstrated that the performances of TME and LNSMI deteriorate in the presence of clustered outliers, whereas the proposed Tyler-type estimator remains stable and provides accurate estimation even in such settings. Based on arguments similar to those in the existing literature, we also proposed a method for selecting the shrinkage parameter of the proposed Tyler-type estimator.
Finally, we outlined several directions for future work. These include the uniqueness of solutions to the minimization problem (7) and the global convergence properties of the proposed method.
@@ Appendix A Reason for focusing on precision estimation in TME and LNSMI
From Figure 3-(b, d, f), we observe that the estimation accuracy of TME and LNSMI decreases monotonically but in a nonlinear manner, whereas that of SCM decreases monotonically and linearly. Based on this observation, we suspected that the difficulty might lie in the part of the procedure related to estimating the precision structure.
@@ Appendix B Number of resampling iterations
$\textnormal{E}[m_{k}]=n\left(\sum_{i=n-k+1}^{n}\frac{1}{i}\right)\text{.}$
@@ Appendix C Generation of the true covariance matrix $\Sigma(c_{1},c_{2})$
     if $\left(\frac{c_{1}/c_{2}}{1-c_{1}/c_{2}-c_{1}}\leq\min\left\{l_{1},\ldots,l_{p-2}\right\}\right)$ then break;
     if $\left(\max\left\{l_{1},\ldots,l_{p-2}\right\}\leq\frac{c_{1}}{1-c_{1}/c_{2}-c_{1}}\right)$ then break;
@@
In Algorithm 1, we first branch according to whether a matrix satisfying the specified largest contribution ratio $c_{1}$ and smallest contribution ratio $c_{1}/c_{2}$ can exist. If such a matrix exists, the contribution ratios are generated in principle from $\textnormal{Dir}_{p-2}(1)$, following the maximum entropy principle. However, depending on the values of $c_{1}$ and $c_{2}$, generating random samples from $\textnormal{Dir}_{p-2}(1)$ may become difficult. In such cases, all contribution ratios from the second to the (p1)(p-1)-th components are set equal.
@@ Appendix E Numerical results under various scenarios -1
Under the numerical experimental settings described in Sections 5.1 and 5.2, we additionally present numerical results for several alternative parameter configurations, following the same format as in Section 5.3.
@@ E.1 Comparison across $\xi$
Figure 6: Plots of the RMSE values of each estimator for different data dimensions $p$, with lines connecting the plotted points. The horizontal axis represents the dimension $p$ $(=2,\ldots,55)$, and the vertical axis represents the RMSE. Figure (a, c) show the results for unclustered outliers, Figure (b, d) correspond to clustered outliers, and Figure (e) represents the case with no outliers, that is, the situation in which all data points are simply generated as random samples from a multivariate normal distribution. The other settings are $N=100$ and $k=10$. In Figure (a, c, e), the curves representing TME, LNSMI, and the proposed method nearly overlap. Furthermore, in Figure (a), the curves representing SCM and $I_{p}$, in Figure (b), those representing TME and LNSMI, and in Figure (e), those representing SCM, TME, LNSMI, and the proposed method also nearly overlap.
@@ E.2 Comparison across $k$
Figure 7: Plots of the RMSE values of each estimator for different data dimensions $p$, with lines connecting the plotted points. The horizontal axis represents the dimension $p$ $(=2,\ldots,55)$, and the vertical axis represents the RMSE. Figure (a, c, e) show the results for unclustered outliers, while Figure (b, d, f) correspond to clustered outliers. The other settings are $N=100$ and $\xi=0.03$. In Figure (a, c, e), the curves representing TME, LNSMI, and the proposed method nearly overlap. Furthermore, in Figure (b, d, f), the curves representing SSCM and the proposed method also nearly overlap.
We now discuss the numerical results. First, as shown in Figures 6 and 7, a substantial degradation in the performance of the TME and the LNSMI is again observed when clustered outliers are present.
Figures 6-(b, d, e) illustrate how the estimator performance varies with respect to the proportion of clustered outliers $\xi$, revealing a characteristic transition behavior. Specifically, in Figure 6-(d), when $\xi=0.03$, a pronounced deterioration in the performance of the TME occurs around dimension $p=33$, whereas in Figure 6-(b), when $\xi=0.1$, a similar phenomenon appears already around $p=10$. These numerical observations are consistent with the fact that the asymptotic breakdown point of the TME lies between $1/(p+1)$ and $1/p$ [9].
Next, Figures 7-(b, d, f) present the estimator performance as a function of the parameter $k$, which controls the severity of clustered outliers. A similar transition behavior can also be observed in this case. Notably, even when $k=1$, that is, in situations where no data points have exceptionally large $\ell_{2}$-norms, a substantial degradation in the performance of both the TME and the LNSMI is still observed.
@@ Appendix F Numerical results under various scenarios -2
Following existing studies [5, 20, 36, 44], we also report numerical results obtained under the $t$-distribution. As in the existing literature, these results indicate that the proposed method successfully attains robustness even when the data follow a $t$-distribution.
Figure 8: Plots of the RMSE values of each estimator for different data dimensions $p$, with lines connecting the plotted points. The horizontal axis represents the dimension $p$ $(=2,\ldots,55)$, and the vertical axis represents the RMSE. Figure (a) shows the results when the data are generated solely from a $t$ distribution with degree of freedom 1, Figure (b) from a $t$ distribution with degree of freedom 3, Figure (c) from a $t$ distribution with degree of freedom 5, and Figure (d) from a $t$ distribution with degree of freedom 10. Note that no outliers are generated separately in this setting. In Figure (a), the curves representing TME, LNSMI, and the proposed method nearly overlap. In Figure (b, c, d), the curves representing TME and the proposed method nearly overlap.

### v2

@@ 6 Proofs
Consider the sequence $\{\mathrm{Y}_{\ell}\}_{\ell\in\mathbb{N}}$ that satisfies $\mathrm{Z}_{\ell}=\mathrm{X}\cup\mathrm{Y}_{\ell}\in\mathcal{E}_{m}(\mathrm{X})$ for $\ell\in\mathbb{N}$. Let
@@
By the compactness of $\mathcal{S}_{+}^{p*}\coloneqq\{X\in\mathcal{S}_{+}^{p};\,\mathrm{Tr}[X]=1\}$ in the Euclidean topology, the sequence $\{\Omega_{\ell}\}_{\ell\in\mathbb{N}}$ admits a convergent subsequence $\{\Omega_{j}\}_{j\in J}$. Hereafter, we shall denote this limit operation by
$\lim_{j\to\infty}\Omega_{j}=\Omega\in\mathcal{S}_{+}^{p*}.$
@@
for $i\in[m]$. Since $\mathbb{S}^{p-1}$ is compact in the Euclidean topology, the product space $\mathcal{S}_{+}^{p*}\times\mathbb{S}^{p-1}$ is also compact in the product topology, and hence there exists an index set $J$ and a point $\bm{u}_{i}\in\mathbb{S}^{p-1}$ such that
$\lim_{j\to\infty}\bm{u}_{i,j}=\bm{u}_{i}.$
Let $\bar{P}_{\Omega}\coloneqq I_{p}-\Omega\Omega^{\dagger}$ denote the orthogonal projection onto $\textnormal{Ker}(\Omega)$. It holds that
@@
Hence, (18) is a necessary condition for $\mathrm{Z}\in\mathcal{E}_{m}(\mathrm{X})$. If $\Omega\bm{x}_{i}\neq\bm{0}$, then
@@
$\lim_{j\to\infty}\psi(\bm{x}_{i}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{x}_{i})\frac{\bm{x}_{i}^{\top}\Omega_{j}^{1/2}\bar{P}_{\Omega}\Omega_{j}^{1/2}\bm{x}_{i}}{\bm{x}_{i}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{x}_{i}}=0.$
@@
$\lim_{j\to\infty}\psi(\bm{y}_{i,j}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{y}_{i,j})\frac{\bm{u}_{i,j}^{\top}\Omega_{j}^{1/2}\bar{P}_{\Omega}\Omega_{j}^{1/2}\bm{u}_{i,j}}{\bm{u}_{i,j}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{u}_{i,j}}=0.$
@@
$\textnormal{rank}(\bar{P}_{\Omega})=\lim_{j\to\infty}\frac{1}{n+m}\sum_{\bm{z}\in\mathrm{Z}_{0,j}}\psi(\bm{z}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{z})\frac{\bm{z}^{\top}\Omega_{j}^{1/2}\bar{P}_{\Omega}\Omega_{j}^{1/2}\bm{z}}{\bm{z}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{z}},$ (19)
@@
$p$ $\geq\lim_{j\to\infty}\frac{1}{n+m}\left\{\sum_{\bm{z}\in\mathrm{Z}_{0,j}}\psi\left(\bm{z}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{z}\right)\frac{\bm{z}^{\top}\Omega_{j}\bm{z}}{\bm{z}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{z}}\right$.
@@
$\geq\lim_{j\to\infty}\frac{1}{n+m}\left\{\sum_{\bm{z}\in\mathrm{Z}_{0,j}}\psi\left(\bm{z}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{z}\right)\frac{\bm{z}^{\top}\Omega_{j}^{1/2}\bar{P}_{\Omega}\Omega_{j}^{1/2}\bm{z}}{\bm{z}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{z}}\right$.
@@
$p$ $\geq\textnormal{rank}(\bar{P}_{\Omega})+\frac{1}{n+m}\sum_{\bm{x}_{i}\in\mathrm{X}\cap\textnormal{Ker}(\Omega)^{c}}\lim_{j\to\infty}\psi\left(\bm{x}_{i}^{\top}(\omega_{j}\Omega_{j}+\alpha I_{p})\bm{x}_{i}\right)\frac{\bm{x}_{i}^{\top}\Omega_{j}\bm{x}_{i}}{\bm{x}_{i}^{\top}\left\{\Omega_{j}+\left(\alpha/\omega_{j}\right)I_{p}\right\}\bm{x}_{i}}.$
Letting $\omega_{j}\to+\infty\,(j\to\infty)$ yields
@@
When $\varepsilon_{m}>1-p/\kappa$, setting $Y=\{\bm{0},\ldots,\bm{0}\}$ yields $\hat{P}_{n+m}(\{\bm{0}\})>1-p/\kappa$, which in turn implies that $Z\notin\mathcal{E}_{m}(\mathrm{X})$, and hence $\hat{V}(\mathrm{Z})$ breaks down, in accordance with Lemma 4.1-(2). When $\varepsilon_{m}<1-p/\kappa-(p-1)/n$, from Lemmas 4.1-(1) and 4.2, we see that $\hat{V}(\mathrm{Z})$ does not breakdown. ∎
@@ 7 Concluding remarks
We further examined the statistical properties of the proposed method from the perspectives of penalized $M$-estimation and Bayesian estimation. Moreover, as summarized in Table 1, we discussed correspondences between the proposed estimator and various existing estimators. Through numerical experiments, we demonstrated that the performance of TME and LNSMI deteriorates in the presence of clustered outliers, whereas the proposed Tyler type estimator remains stable and provides accurate estimation even in such settings. Based on arguments similar to those in the existing literature, we also proposed a method for selecting the shrinkage parameter of the proposed Tyler type estimator.
Finally, we outlined several directions for future work. These include, but are not limited to, the uniqueness of solutions to the minimization problem (7) and the global convergence properties of the proposed estimator.
@@ Appendix A Reason for focusing on precision estimation in TME and LNSMI
From Figure 3-(b, d, f), we observe that the estimation accuracy of TME and LNSMI decreases monotonically but in a nonlinear manner, whereas that of SCM decreases monotonically and linearly. Based on this observation, we suspected that the difficulty lies in the part of the procedure related to estimating the precision structure.
@@ Appendix B Number of resampling iterations
$\mathrm{E}[m_{k}]=N\left(\sum_{i=N-k+1}^{N}\frac{1}{i}\right)\text{.}$
@@ Appendix C Generation of the true covariance matrix $\Sigma(c_{1},c_{2})$
     if $\left(\frac{c_{1}/c_{2}}{1-c_{1}/c_{2}-c_{1}}\leq\min\left\{l_{1},\ldots,l_{p-2}\right\}\right)$ and $\left(\max\left\{l_{1},\ldots,l_{p-2}\right\}\leq\frac{c_{1}}{1-c_{1}/c_{2}-c_{1}}\right)$ then break;
@@
In Algorithm 1, we first branch according to whether a matrix satisfying the specified largest contribution ratio $c_{1}$ and smallest contribution ratio $c_{1}/c_{2}$ can exist. If such a matrix exists, the contribution ratios are generated from $\textnormal{Dir}_{p-2}(1)$ in principle. However, depending on the values of $c_{1}$ and $c_{2}$, generating random samples from $\textnormal{Dir}_{p-2}(1)$ may become difficult. In such cases, all contribution ratios from the second to the (p1)(p-1)-th components are set equal.
@@ Appendix D Table of RMSE values corresponding to Figure 3
Table 2 reports the RMSE values corresponding to Figure 3.
@@ Appendix E Numerical results under various scenarios -1
Under the numerical settings described in Sections 5.1 and 5.2, we additionally present numerical results for several alternative parameter configurations, following the same format as in Section 5.3.
@@ E.1 Comparison across $\xi$
Figure 6: Plots of the RMSE values of each estimator for different data dimensions $p$, with lines connecting the plotted points. The horizontal axis represents the dimension $p$ $(=2,\ldots,55)$, and the vertical axis represents the RMSE. Figure (a, c) show the results for unclustered outliers, Figure (b, d) correspond to clustered outliers, and Figure (e) represents the case with no outliers, that is, the situation in which all data points are simply generated as random samples from a multivariate normal distribution. The other settings are $N=100$ and $k=10$. In Figure (a, c, e), the curves representing TME, LNSMI, and the proposed Tyler type estimator nearly overlap. Furthermore, in Figure (a), the curves representing SCM and $I_{p}$, in Figure (b), those representing TME and LNSMI, and in Figure (e), those representing SCM, TME, LNSMI, and the proposed Tyler type estimator also nearly overlap.
We now discuss the numerical results. First, as shown in Figure 6, a substantial degradation in the performance of the TME and the LNSMI is again observed when clustered outliers are present.
Figure 6-(b, d, e) illustrates how the estimator performance varies with respect to the proportion of clustered outliers $\xi$, revealing a characteristic transition behavior. Specifically, in Figure 6-(d), when $\xi=0.03$, a pronounced deterioration in the performance of the TME occurs around dimension $p=33$, whereas in Figure 6-(b), when $\xi=0.1$, a similar phenomenon appears already around $p=10$. These numerical observations are consistent with the fact that the asymptotic breakdown point of the TME lies between $1/(p+1)$ and $1/p$ [8].
@@ E.2 Comparison across $k$
Figure 7: Plots of the RMSE values of each estimator for different data dimensions $p$, with lines connecting the plotted points. The horizontal axis represents the dimension $p$ $(=2,\ldots,55)$, and the vertical axis represents the RMSE. Figure (a, c, e) show the results for unclustered outliers, while Figure (b, d, f) correspond to clustered outliers. The other settings are $N=100$ and $\xi=0.03$. In Figure (a, c, e), the curves representing TME, LNSMI, and the proposed Tyler type estimator nearly overlap. Furthermore, in Figure (b, d, f), the curves representing SSCM and the proposed Tyler type estimator also nearly overlap.
First, as shown in Figure 7, a substantial degradation in the performance of the TME and the LNSMI is again observed when clustered outliers are present.
Next, Figure 7-(b, d, f) presents the estimator performance as a function of the parameter $k$, which controls the severity of clustered outliers. A similar transition behavior can also be observed in this case. Notably, even when $k=1$, that is, in situations where no data points have exceptionally large $\ell_{2}$-norms, a substantial degradation in the performance of both the TME and the LNSMI is still observed.
@@ Appendix F Numerical results under various scenarios -2
Following existing studies [5, 19, 35, 43], we also report numerical results obtained under the $t$-distribution. As in the existing literature, these results indicate that the proposed Tyler type estimator successfully attains robustness even when the data follow a $t$-distribution.
In this section, the data, including outliers, are generated from a $t$-distribution. This setting corresponds to the case of unclustered outliers, and none of the estimators exhibit breakdown behavior. The curves representing the performance of TME, LNSMI, SSCM, and the proposed Tyler type estimator nearly overlap.
Figure 8: Plots of the RMSE values of each estimator for different data dimensions $p$, with lines connecting the plotted points. The horizontal axis represents the dimension $p$ $(=2,\ldots,55)$, and the vertical axis represents the RMSE. Figure (a) shows the results when the data are generated solely from a $t$ distribution with degree of freedom 11, Figure (b) from a $t$ distribution with degree of freedom 3, Figure (c) from a $t$ distribution with degree of freedom 5, and Figure (d) from a $t$ distribution with degree of freedom 10. Note that no outliers are generated separately in this setting. In Figure (a), the curves representing TME, LNSMI, and the proposed Tyler type estimator nearly overlap. In Figure (b, c, d), the curves representing TME and the proposed Tyler type estimator nearly overlap.